OpenStudy (anonymous):

Choose the correct simplification of (6x6y12z3)0(2x)3.

OpenStudy (anonymous):

(6x^6y^12z^3)^0(2x)^3

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@amistre64

OpenStudy (mertsj):

\[(6x^6y ^{12}z^3)^0(2x)^3\]

OpenStudy (mertsj):

Do I have it right?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

We know that anything raised to the power of 0 is 1 so the first part is 1. So now we have: \[(2x)^3=2^3x^3=8x^3\]